OpenStudy (anonymous):

how to solve this problem??? If a ball is dropped from the top of a 256-ft building, then the formula h(t)= 256-16t^2expresses its height h(t) in feet as a function of the time t in seconds. a. Find h(2), the height of the ball 2 seconds after it is dropped. b. Find h(4) -topic about function

OpenStudy (anonymous):

help please .. :)

OpenStudy (anonymous):

h(t)=256-16y^2 a. h(2)=256-16(2^2) h*2=256-16*4 h*2=256-64 h*2=192 divide both sides by 2 h=96 b. h(4)=256-16(4^2) h*4=256-16*16 h*4=256-256 h*4=0 divide both sides by 4 h=0

OpenStudy (ybarrap):

$$ h(t)=256-16y^2\\ a.\\ h(2)=256-16(2^2)\\ h(2)=256-16*4\\ h(2)=256-64\\ h(2)=192\\ \cancel{\text{divide both sides by 2 h=96}} \\ \\ b.\\ h(4)=256-16(4^2)\\ h(4)=256-16*16\\ h(4)=256-256\\ h(4)=0 \\ \cancel{\text{divide both sides by 4 h=0}} $$